OpenStudy (readergirl12):

What is the best way to classify each equation? Column A Column B A. 1. 8x + 24=2(4x+12) identity 2. 5x+8-x=2(2x+8) B. 3. 7(3x-2)=20x-13 contradiction 4. 3x+2(x-10)=5(x-4) C. neither

OpenStudy (readergirl12):

Help.. v-v

OpenStudy (anonymous):

8x+24= 2(4x+12)=2*4x+12*2=8x+24

OpenStudy (anonymous):

so we see that it if identity

OpenStudy (anonymous):

5x+8-x=2(2x+8) but 2(2x+8)=2*2x+2*8 which gives 4x+16

OpenStudy (anonymous):

i gather that contradiction mean... The method of proof by contradiction is to assume that a statement is not true and then to show that that assumption leads to a contradiction. In the case of trying to prove this is equivalent to assuming that That is, to assume that is true and is false.

OpenStudy (anonymous):

so im not sure if that 1 is contradition or neither

OpenStudy (anonymous):

7(3x-2)=20x-13 7*3x-7*2= 21x-14 im thinking this one shd be contradiction

OpenStudy (anonymous):

3x+2(x-10)=5(x-4) 3x+2(x-10)= 3x+2*x-2*10=3x+2x-20=5x-20 5(x-4)=5*x-5*4=5x-20 so identify